OpenStudy (anonymous):

16 is what percent of 80 how do i do this???

OpenStudy (anonymous):

is over of % over 100

OpenStudy (anonymous):

16 / 100

jhonyy9 (jhonyy9):

so 80 is 100% 16 --- x --------- x = 16*100/80 = ? %

OpenStudy (anonymous):

i thought it was the % * the number?

OpenStudy (anonymous):

ok i got 2 different answers is there any answers to this question @nisa ?

OpenStudy (anonymous):

yes its a college placement test practice thing it says the answer is 20 %

jhonyy9 (jhonyy9):

so than check it my answer

OpenStudy (anonymous):

that is right thats what i got

OpenStudy (anonymous):

@jhonyy9 yeah you are right :)

OpenStudy (anonymous):

is/of and %/100 16/x 80/100 cross multiply 16 x 100 and 80x X 1600 and 80x divide 1600/80 =20%

jhonyy9 (jhonyy9):

so remember ALLWAYS you need using the rule of 3 simple and so will get the right answer easy

OpenStudy (anonymous):

so the first number times a hundred divided by whatever number is given?

jhonyy9 (jhonyy9):

give an different example

jhonyy9 (jhonyy9):

so 25 is how many percent of 100 ?

jhonyy9 (jhonyy9):

how you write it now for you can giving the right answer ?

OpenStudy (anonymous):

25*100/100?

jhonyy9 (jhonyy9):

100 is 100% 25 is --- x % ------------- x= 25*100/100 = 25

jhonyy9 (jhonyy9):

or you need a different exemple again ?

OpenStudy (anonymous):

i think i got how to do it :) thanks alot :)

jhonyy9 (jhonyy9):

3 is how many pecent of 9 ?

OpenStudy (anonymous):

3*100/9 33.33 repeating

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so this mean that you have learned it sure now

jhonyy9 (jhonyy9):

yes ? this is easy

OpenStudy (anonymous):

yes i have thanks so much :)

OpenStudy (anonymous):

it seems really easy now

jhonyy9 (jhonyy9):

yw was my pleasure that i like learning people math good luck bye

OpenStudy (anonymous):

good luck to u too i appreciate it :) byee

OpenStudy (anonymous):

thanks everybody :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

:)